# Combo: Co-speech holistic 3D human motion generation and efficient customizable adaptation in harmony

### v1

@@ 1 Introduction
In light of the above, we examine the holistic human motion generation task from a principled perspective and advocate two novel designs tailored for improving data-driven generative model. Our key insight is to alleviate modeling complexity in both input and output end, leading to a harmonious but also adaptive generative network. For the input end, instead of training an encoder that consumes all at once, we pre-train a model in a relatively clear setting, and defer the injection of important conditions to the fine-tuning stage; For the output end, we propose a divide-and-unite strategy to address the trade-off between learning accurate facial and body movements and guaranteeing coherence. In the following, we motivate and explain the above designs in more detail.
@@
Regarding the multiple input end, our generative model is expected to take speeches as generation guidance, which contain factors such as contents, rhythms, semantics, and so on. It is worth noting that the same speech guidance can lead to different motion sequences with respect to varying conditions (e.g., identity and emotion), complicating the learning process. The significant contributions to high-quality dataset construction [5, 3], containing rich annotations as well as great variability in identity and emotions, offer good opportunities for learning a comprehensive model. However, we argue that learning a comprehensive model lacks flexibility in further adaptation. In fact, generative models have been widely desired to be controllable and customizable [6, 7, 8, 9]. Regarding holistic human motion generation, one would naturally expect a flexible, adaptive digital avatar with respect to changes in both short-term emotional states (emotion) and long-term personality (identity). Unfortunately, recent efforts [1, 3, 2, 10, 11] require an extensive training or full-parameter fine-tuning process to accommodate to the emotional data or newly introduced identities, a limitation that is particularly exacerbated in VQ-VAE-based methods [1, 3], as shown in Fig. 1(b).
@@
Motivated by the above observations, we take a pre-train-and-fine-tune approach. During pre-training, we train a model with data from a fixed identity in neutral emotion. The injection of specific conditions, such as alternative identities and/or emotions, is deferred to the fine-tuning stage. Our key design in this part is a novel plug-and-play adapter for Parameter-Efficient Fine-Tuning (PEFT) [12], X-Adapter, which is tailored to our pipeline and allows for effective and efficient fine-tuning on emotion transfer and personalized generation. Our treatment on input end admits several advantages: 1) It alleviates the learning burden during pre-train stage, therefore improving the performance; 2) Built upon the advanced pre-training, the fine-tuning phase can achieve superior performance with minimal computational cost; 3) As a by-product, it further enables versatile editing during inference benefiting from flexible X cues, which is a capability that previous approaches [1, 2, 3] could not attain. For instance, we can employ multi-modal conditions within the CLIP domain [13] to indicate the short-term emotion, which supports flexible and generalizable zero-shot editing. We can also define identity codes as several interpretable statistics of SMPLX [14] to depict long-term personality.
@@ 2.2 Parameter-Efficient Fine-Tuning
Unlike conventional fine-tuning, which updates all parameters, Parameter-Efficient Fine-Tuning (PEFT) has shown impressive results across various tasks by updating only a subset of parameters. The three primary methods of PEFT are Adapter [43, 44], prefix-tuning [45, 46], and LoRA [47, 48]. In the realm of NLP, the concept of adapters was first introduced by the Serial Adapter [49] model. This model enhances each Transformer block by incorporating two additional adapter modules, one after the self-attention layer and the other after the feed-forward neural network (FFN) layer. Parallel Adapter [50], on the other hand, restructures these sequential layers into a parallel side network that operates in tandem with each Transformer sublayer. Our X-Adapter is largely based on this concept. With the growing interest in diffusion models [29, 51], recent studies have applied PEFT to refine pre-trained diffusion models for specialized tasks. We introduce two key applications, including integrating additional input modalities and customizing content generation. Specifically, GLIGEN [52] integrates new trainable gated Transformer layers without altering the original model’s weights, while ControlNet [53] fine-tunes a separate encoding layer copy from Stable Diffusion [51] while locks its pre-trained parameter weights. T2I-Adapter [54] adds a lightweight model to synchronize external controls with the model’s internal processes. For customization, Textual Inversion [55] identifies pseudo-words for concept representation, and IP-Adapter [56] inserts an image-focused attention layer for feature enhancement. Notably, EAT [57] employs a pre-train-and-fine-tune approach similar to ours, but they extend a trainable encoder branch for fine-tuning directly, which is neither as efficient nor as effective as PEFT. In this work, we integrate PEFT into the domain of holistic 3D human motion generation for flexible adaptation, which is the first attempt in this field.
@@ 3.2 DU-Trans
where []delimited-[][$\cdot$][ ⋅ ] means concatenation, $l$ is the number of encoder layers, and PE is the positional embedding. The body branch, on the other hand, receives semantic information rather than phonetic content:
@@
3) A Bi-Flow layer built upon the cross-attention mechanism is introduced to preliminarily model the relationship between the two branches, capturing holistic dynamic priors to enhance the performance of each component. We take face to body data-flow for example, the query $\boldsymbol{Q}_{B}$ is extracted by linear projection from $\boldsymbol{f}^{B}_{i}$ ($i$ is the layer index), and the key and value $\boldsymbol{K}_{F}$, $\boldsymbol{V}_{F}$ is extracted from $\boldsymbol{f}^{F}_{i}$ in the same way. To obtain the updated body feature $\tilde{\boldsymbol{f}}^{B}_{i}$,
@@
4) Our architecture consists of three output heads: Two output heads are used to predict independent coefficients for face and body under auxiliary supervisions (c.f. Loss Functions paragraph below); One output head is for holistic generation. By summing the face features $\boldsymbol{f}_{l}^{F}$ and body features $\boldsymbol{f}_{l}^{B}$ output by the two encoders and feeding the sum into the single decoder $\boldsymbol{\Phi}^{FB}$ to implicitly model their interrelations, During training, all output heads are optimized with regarding training loss, while during inference only the last head is activated to deliver coordinated holistic human motions. Formally,
@@
Loss Functions. During the training phase, our model outputs three predicted parameters: facial expression $\hat{\boldsymbol{F}}^{1:N}\in\mathbb{R}^{N\times 100}$ and body gestures $\hat{\boldsymbol{B}}^{1:N}\in\mathbb{R}^{N\times 165}$ output from two encoder, and the combined one $\hat{\boldsymbol{H}}^{1:N}\in\mathbb{R}^{N\times 265}$ output from a single decoder. Each is supervised by two loss functions, simple loss $\mathcal{L}_{\text{simple }}$ following the Eq. 3 and velocity loss $\mathcal{L}_{\text{vel}}$. Formally, we take the holistic body as an example:
@@ 3.3 X-Adapter
2) We insert a modulation layer $\boldsymbol{\mathcal{M}}$ before the down-projection phase to seamlessly infuse task-specific conditions X ($\boldsymbol{z}_{s}$ or $\boldsymbol{z}_{id}$) into the adapter module. Striking a balance between performance and parameter efficiency, this modulation is elegantly achieved through the use of addition alone:
@@ 4.1.2 Implementation Details
We implement our network using PyTorch. The basic DU-Trans contains seven encoder layers and one decoder layer. The hidden dimension of all transformer layers is 512. The bottleneck dimension of X-Adapter is set to 128. Our diffusion model employs a cosine noise schedule, with diffusion steps set to 1000 for training and inference. During pretraining, the learning rate is set to 1e-4 using the ADAM optimizer with $[\beta_{1},\beta_{2}]=[0.9,0.99]$ and adjusted to 1e-3 for fine-tuning experiments. The batch size is consistently set to 128. We train our pretraining model on a single NVIDIA A100 GPU for one day, completing 100,000 iterations. For fine-tuning, we perform 5,000 iterations within one hour. The length of the input motion sequences is 600.
4.1.3 Metrics
We assess the quality of the generated motion from three aspects. For the whole body, we utilize FMD [2] to measure the difference between the distributions of the generated holistic motion and ground truth in feature space. It can indicate the overall quality of the holistic motions and the coordination among different body parts. For body gestures, we use FGD [69] to measure the distribution difference between generated gesture and ground truth. BC [70] quantifies the alignment between the rhythm of the generated gesture and the beat of the audio. DIV [36] is a metric for measuring the variations of the synthesized gesture. For face expression, we employ two reconstruction metrics. The vertex MSE [28] is calculated to determine the positional distance, and the vertex L1 difference LVD [1] is used to assess the discrepancy between the ground truth and the generated facial vertices. Besides, since this task lacks clear ground, human objective evaluation is another main criterion in our method. We conduct an extensive user study to rate the generated motions by different methods in terms of coordination, coherence, and synchronization.
Figure 5: Qualitative Comparison with TalkSHOW and EMAGE on SHOW Dataset. At the left part, we visualize the lower body for all methods.
4.1.4 Baselines
For BEAT2 dataset [3], we compare our method with representative SOTA approaches in talking head generation: FaceFormer [21] and CodeTalker [28], both of which are tailored for speech-driven 3D facial animation. Additionally, we evaluate against body gesture generation methods, CaMN [5] which introduce the BEAT dataset constructed by using a commercial motion capture system, and DSG [37], a novel diffusion-based co-speech generation method. We reproduce their results in the standardized format of BEAT2 for the face and body respectively. Furthermore, we incorporate assessments of three recent holistic pipelines: the Habibie et al. [35] and TalkSHOW [1] are retrained in standardized format either, and EMAGE [3] which directly uses their officially released weights for evaluation. On the SHOW dataset [1], all methods require retraining, including TalkSHOW, as its official weights are general across all identities, whereas other comparison methods like EMAGE are trained for single identity only. To ensure a fair comparison, we retrain TalkShow for each identity. We do not include DiffSHEG [2] and ProbTalk [4] in our comparison since we failed to obtain correct results based on their released codes when retraining them. A more detailed comparison with them will be provided later once the issues are resolved.
Table I: Quantitative comparison with SOTA methods on BEAT2 dataset. The $\downarrow$” means the lower, the better, and vice versa. Bold and underline represent optimal and suboptimal results. The * indicates training from scratch (pretraining), while the $\dagger$ signifies fine-tuning the emotional model from the neutral pre-trained one. For simplicity, we report MSE×$\times 10^{-8}$ and LVD×$\times 10^{-5}$ as EMAGE.
@@ 4.2.1 Quantitative Results
We first consider a clear training set, Neutral video clips of Scott from BEAT2. We show that our method, trained from scratch (indicated by *), can achieve superior performance than the competing methods. As shown in the top part of Tab. I, we begin by analyzing the synchronization of facial expressions, our method significantly outperforms the baselines in both MSE and LVD by a large margin, applicable to both specialized talking head generation methods [28, 21] and other holistic motion generation methods [41, 1, 3]. Then, we shift our focus to the quality of gestures, where previous co-speech methods [5, 37] struggle to match our results. Moreover, our method also outperforms TalkSHOW and EMAGE in terms of FGD and BC, with FGD being particularly notable, which indicates that the distribution of our results is the closest to the ground truth. Note that our method’s DIV metric is slightly lower than that of TalkSHOW and EMAGE, which can be attributed to the fact that the comparative methods include some sudden and exaggerated meaningless movements, leading to a higher score. Finally, we evaluate the motion produced by holistic methods from an comprehensive viewpoint. Our approach shows remarkable advancements in FMD, indicating that it excels in generating synchronized and coordinated holistic body motions. Overall, thanks to the divide and unite mechanism in DU-Trans, our method not only demonstrates superior performance in the metrics of individual components but also clearly leads to holistic metrics. Similar observations could also be concluded from the quantitative results on SHOW dataset, as shown in Tab. II. Notably, the solid pre-training performance brought by this powerful basic architecture will benefit the subsequent fine-tuning stage. For more details, please refer to Sec. 4.3.
@@
CaMN [5] 3.365 2.199 0.7998 10.13 - -
DSG [37] 3.462 2.404 0.8295 10.04 - -
Habibie et al. [35] 3.851 2.679 0.8510 8.055 145.1 47.11
TalkSHOW [1] 3.478 2.462 0.8449 10.29 139.3 44.81
EMAGE [3] 3.380 2.255 0.8585 12.40 136.4 42.74
Ours* 3.142 2.067 0.8667 10.36 133.5 38.21
@@ 4.2.3 User Study
It is widely acknowledged that assessing the quality of generative tasks is inherently subjective. Despite that we have evaluated with multiple quantitative metrics, there remains a significant gap between such and human visual perception. To this end, we conduct a user study to compare our framework with two recent baselines, TalkSHOW and EMAGE.
Specifically, we randomly select 100 synthesized samples of Scott, including various neutral and emotional clips. Then we recruit 100 subjects with diverse backgrounds to select the most preferred motions in terms of the whole-body coordination, holistic coherence, and synchronization of the expression and gesture with the audio for the shuffled visual results. Besides, in order to help the subjects get used to our questionnaire, we discard the answers of the first three samples and append them in the end to allow participants to get used to the task (test on 103 clips and take the last 100 answers). The results are shown in Tab. III, our generated holistic motions are dominantly preferred on all three metrics over the competing baselines, especially on coordination and synchronization, which aligns with the superior performance observed in FMD, FGD, and MSE as illustrated in the Tab. I and Tab. II. Thus it can be concluded that DU-Trans extracts the distinctive features of both the face and body while enhancing the harmony between them. Overall, our approach is capable of generating more coordinated, coherent, and synchronized motions that humans prefer.
4.3 Verification on X-Adapter
In Sec. 4.3.1, we first analyze the effectiveness of the X-adapter in emotion transfer during training, and then further explore the two benefits it brings: multi-modal editing and unseen emotion editing during inference. These experiments are conducted on Scott data. Then, in Sec. 4.3.2, we analyze the effectiveness of the X-adapter in identity transfer and further validate cross-identity transfer under various conditions. These experiments are conducted on Neutral data. Finally, in Sec. 4.3.3, we provide a detailed analysis of overall tuning efficiency.
Table III: Results of the user study. This shows user preference percentage for coordination, coherence, and synchronization. Given that all these methods generate temporally coherent motions, their coherence results are comparably similar. Our methods are mainly preferred for coordination and synchronization.
Method Coordinated $\uparrow$ Coherent $\uparrow$ Synchronized $\uparrow$
TalkSHOW [1] 0.16 0.30 0.21
EMAGE [3] 0.38 0.34 0.35
Ours 0.46 0.36 0.44
4.3.1 Analysis on Emotion Transfer

### v2

@@ 1 Introduction
In light of the above, we examine the holistic human motion generation task from a principled perspective and advocate two novel designs tailored for improving data-driven generative model. Our key insight is to alleviate modeling complexity in both input and output end, leading to a harmonious but also adaptive generative network. For the input end, instead of training an encoder that consumes all at once, we pre-train a model in a relatively clear setting, and defer the injection of important conditions to the fine-tuning stage; For the output end, we propose a divide-and-unite strategy to address the trade-off between learning accurate facial and body movements and guaranteeing coherence. In below, we motivate and explain the above designs in more detail.
@@
Regarding the multiple input end, our generative model is expected to take speeches as generation guidance, which contain factors such as contents, rhythms, semantics, etc. It is worth noting that the same speech guidance can lead to different motion sequences with respect to varying conditions (e.g., identity and emotion), complicating the learning process. The significant contributions to high-quality dataset construction [5, 3], containing rich annotations as well as great variability in identity and emotions, offer good opportunities for learning a comprehensive model. However, we argue that learning a comprehensive model lacks flexibility in further adaptation. In fact, generative models have been widely desired to be controllable and customizable [6, 7, 8, 9]. Regarding holistic human motion generation, one would naturally expect a flexible, adaptive digital avatar with respect to changes in both short-term emotional states (emotion) and long-term personality (identity). Unfortunately, recent efforts [1, 3, 2, 10, 11] require an extensive training or full-parameter fine-tuning process to accommodate to the emotional data or newly introduced identities, a limitation that is particularly exacerbated in VQ-VAE-based methods [1, 3], as shown in Fig. 1(b).
@@
Motivated by the above observations, we take a pre-train-and-fine-tune approach. During pre-training, we train a model with data from a fixed identity in neutral emotion. The injection of specific conditions, such as alternative identities and/or emotions, is deferred to the fine-tuning stage. Consequently, this technique overcomes the complexity of training a universal model with various conditions by a two-stage scheme, i.e., first train on a simple condition to establish a basic model, and then perform fast fine-tuning to adapt this model to different scenarios. Our key design in this part is a novel plug-and-play adapter for Parameter-Efficient Fine-Tuning (PEFT) [12], X-Adapter, which is tailored to our pipeline and allows for effective and efficient fine-tuning on emotion transfer and personalized generation. Specifically, our treatment on input end admits several advantages: 1) It eliminates complex conditions by training a simple basic model, which can be trained easier and more through for improved performance; 2) Leveraging the advanced pre-training, the fast fine-tuning phase requires only minimal parameter adjustments to guide the model in the desired conditional direction while maintaining excellent performance; 3) As a by-product, it further enables versatile editing during inference benefiting from flexible X cues, which is a capability that previous approaches [1, 2, 3] could not attain. For emotion, we can employ multi-modal conditions within the CLIP domain [13] to indicate the short-term emotion, which supports flexible and generalizable zero-shot editing. For identity, we define identity codes as several interpretable statistics derived from the face and body parameters of SMPLX [14] to depict long-term personality.
@@ 2.2 Parameter-Efficient Fine-Tuning
Unlike conventional fine-tuning, which updates all parameters, Parameter-Efficient Fine-Tuning (PEFT) has shown impressive results across various tasks by updating only a subset of parameters, with the three primary methods being Adapter, Prefix-tuning, and LoRA. First, the concept of adapters is first introduced by the Serial Adapter [43] model. This model enhances each Transformer block by incorporating two additional adapter modules, one after the self-attention layer and the other after the feed-forward neural network (FFN) layer. Parallel Adapter [44], on the other hand, restructures these sequential layers into a parallel side network that operates in tandem with each Transformer sublayer. Other works, such as CIAT [45] and CoDA [46] are also adopts a parallel adapter. Second, Prefix-tuning [47] introduces tunable vectors that are prepended to keys and values across all attention layers. Concurrent work p-tuning [48] and prompt-tuning [49] apply learnable vectors only at the initial word embedding layer rather than all layers to enhance training and inference efficiency. Although these concepts can be prone to slow convergence, they have been employed for various downstream tasks [50, 51]. Third, LoRA [52] adopts a low-rank approximation to update weight matrices in the attention. Several subsequent studies [53, 54, 55] aim to improve LoRA’s performance in various aspects. Our X-Adapter is largely based on the first choice and follows the parallel design. Notably, EAT [56] employs a pre-train-and-fine-tune approach similar to ours, but they extend a trainable encoder branch for fine-tuning directly, which is neither as efficient nor as effective as PEFT.
In this work, we integrate PEFT into the domain of holistic 3D human motion generation for fast and flexible adaptation, which is the first attempt in this field. From an application perspective, holistic 3D human motion finds applications across various scenarios. However, currently there is no single universal model capable of handling all demands. A fallback approach is to offer only some customized services, which, however, lacks flexibility and limits the scope of application. Our motivation is to have a powerful basic model that can cover most demands, and combine it with PEFT for fast fine-tuning. This allows the basic model to quickly adapt to various fine-tuning data within a short period, thereby covering a wide range of scenarios. We believe that the combination of pretraining and fine-tuning is a viable approach to advancing the practical application of motion generation. More methodological and experimental details are presented in the Secs. 3-4.
@@ 3.2 DU-Trans
where $\cdot$] means concatenation, $l$ is the number of encoder layers, and PE is the positional embedding. The body branch, on the other hand, receives semantic information rather than phonetic content:
@@
3) A Bi-Flow layer built upon the cross-attention mechanism [66] is introduced to preliminarily model the relationship between the two branches, capturing holistic dynamic priors to enhance the performance of each component. We take face to body data-flow for example, the query $\boldsymbol{Q}_{B}$ is extracted by linear projection from $\boldsymbol{f}^{B}_{i}$ (ii is the layer index), and the key and value $\boldsymbol{K}_{F}$, $\boldsymbol{V}_{F}$ is extracted from $\boldsymbol{f}^{F}_{i}$ in the same way. To obtain the updated body feature $\tilde{\boldsymbol{f}}^{B}_{i}$,
@@
4) Our architecture consists of three output heads: Two output heads are used to predict independent coefficients for face and body under auxiliary supervisions (c.f. Loss Functions paragraph below); One output head is for holistic generation. By summing the face features $\boldsymbol{f}_{l}^{F}$ and body features $\boldsymbol{f}_{l}^{B}$ output by the two encoders and feeding the sum into the single decoder $\boldsymbol{\Phi}^{FB}$ to implicitly model their interrelations. During training, all output heads are optimized with regarding training loss, while during inference only the last head is activated to deliver coordinated holistic human motions. Formally,
@@
[…]
@@ 3.3 X-Adapter
[…]
@@ 4.1.2 Implementation Details
We implement our network using PyTorch. The basic DU-Trans contains seven encoder layers and one decoder layer. The hidden dimension of all transformer layers is 512. The bottleneck dimension of X-Adapter is set to 128. Our diffusion model employs a cosine noise schedule, with diffusion steps set to 1000 for training and inference. During pretraining, the learning rate is set to 1e-4 using the ADAM optimizer with $\beta_{1},\beta_{2}]=[0.9,0.99]$ and adjusted to 1e-3 for fine-tuning experiments. The batch size is consistently set to 128. We train our pretraining model on a single NVIDIA A100 GPU for one day, completing 100,000 iterations. For fine-tuning, we perform 5,000 iterations within one hour. The length of the input motion sequences is 600.
[…]
@@
BEAT2 (Nidal) Neutral TalkSHOW [1] 0.706 0.256 0.6213 5.693 1.825 4.118
EMAGE [3] 0.688 0.235 0.6272 6.057 1.983 4.021
Ours* 0.651 0.228 0.6391 6.062 1.685 3.768
Emotional TalkSHOW [1] 0.817 0.322 0.6284 6.796 2.897 4.821
EMAGE [3] 0.801 0.314 0.6405 7.308 2.896 4.705
[…]
4.1.3 Metrics
We assess the quality of the generated motion from three aspects. For the whole body, we utilize FMD [2] to measure the difference between the distributions of the generated holistic motion and ground truth in feature space. It can indicate the overall quality of the holistic motions and the coordination among different body parts. For body gestures, we use FGD [69] to measure the distribution difference between generated gesture and ground truth. For the above two metrics, we use a Skeleton CNN based encoder and a Full CNN-based decoder as the autoencoder pretrained network following [3]. The network is pretrained on both BEAT2 and SHOW datasets. BC [70] quantifies the alignment between the rhythm of the generated gesture and the beat of the audio. DIV [36] is a metric for measuring the variations of the synthesized gesture. For face expression, we employ two reconstruction metrics. The vertex MSE [28] is calculated to determine the positional distance, and the vertex L1 difference LVD [1] is used to assess the discrepancy between the ground truth and the generated facial vertices. Besides, since this task lacks clear ground, human objective evaluation is another main criterion in our method. We conduct an extensive user study to rate the generated motions by different methods in terms of coordination, coherence, and synchronization.
Figure 5: Qualitative comparison with TalkSHOW and ProbTalk on SHOW dataset. We utilize official checkpoints of competing models and follow the official rendering format for fair comparisons.
4.1.4 Baselines
For BEAT2 dataset [3], we compare our method with representative SOTA approaches in talking head generation: FaceFormer [21] and CodeTalker [28], both of which are tailored for speech-driven 3D facial animation. Additionally, we evaluate against body gesture generation methods, CaMN [5] which introduce the BEAT dataset constructed by using a commercial motion capture system, and DSG [37], a novel diffusion-based co-speech generation method. We reproduce their results in the standardized format of BEAT2 for the face and body respectively. Furthermore, we incorporate assessments of three recent holistic pipelines: the Habibie et al. [35] and TalkSHOW [1] are retrained in standardized format either, and EMAGE [3] which directly uses their officially released weights for evaluation. On the SHOW dataset [1], we replace EMAGE with the recent ProbTalk [4] because it offers an official checkpoint. Additionally, TalkSHOW also utilizes the officially released weights. The motion rendering follows the official SHOW format, omitting the lower body in a seated position to ensure a fair comparison. We do not include DiffSHEG [2] in our comparison since we fail to correctly visualize the SHOW results.
@@ 4.2.1 Quantitative Results
We first consider a clear training set, Neutral video clips of Scott from BEAT2. We show that our method, trained from scratch (indicated by *), can achieve superior performance than the competing methods. As shown in the BEAT2(Scott)-Neutral part of Tab. I, we begin by analyzing the synchronization of facial expressions, our method significantly outperforms the baselines in both MSE and LVD by a large margin, applicable to both specialized talking head generation methods [28, 21] and other holistic motion generation methods [41, 1, 3]. Then, we shift our focus to the quality of gestures, where previous co-speech methods [5, 37] struggle to match our results. Moreover, our method also outperforms TalkSHOW and EMAGE in terms of FGD and BC, with FGD being particularly notable, which indicates that the distribution of our results is the closest to the ground truth. Note that our method’s DIV metric is slightly lower than that of TalkSHOW and EMAGE, which can be attributed to the fact that the comparative methods include some sudden and exaggerated meaningless movements, leading to a higher score. Finally, we evaluate the motion produced by holistic methods from an comprehensive viewpoint. Our approach shows remarkable advancements in FMD, indicating that it excels in generating synchronized and coordinated holistic body motions. Besides, we present quantitative evaluations on Beat2-Nidal, an identity that is distinctly different from Scott. Our method consistently verifies superior performance compared to the recent TalkSHOW and EMAGE, as shown in the BEAT2(Nidal) part of Tab. I.
Overall, thanks to the divide and unite mechanism in DU-Trans, our method not only demonstrates superior performance in the metrics of individual components but also clearly leads to holistic metrics. Similar observations could also be concluded from the quantitative results on SHOW dataset, as shown in Tab. II. Notably, the solid pre-training performance brought by this powerful basic architecture will benefit the subsequent fine-tuning stage. For more details, please refer to Sec. 4.3.
@@
CaMN [5] 3.187 2.128 0.8035 9.98 - -
DSG [37] 3.253 2.349 0.8266 10.12 - -
Habibie et al. [35] 3.440 2.643 0.8415 8.007 145.1 47.11
TalkSHOW [1] 3.038 2.107 0.8590 10.25 139.3 44.81
ProbTalk [4] 3.274 2.139 0.8505 10.04 151.7 49.66
Ours* 2.961 2.002 0.8619 10.74 136.0 40.92
@@ 4.2.3 User Study
It is widely acknowledged that assessing the quality of generative tasks is inherently subjective. Despite that we have evaluated with multiple quantitative metrics, there remains a significant gap between such and human visual perception [71, 72]. To this end, we conduct rigorous user studies to compare our framework with three baselines: mocap recordings (GT), TalkSHOW [1], and EMAGE [3] on BEAT2 dataset.
Specifically, we utilize the official code from EMAGE to visualize the motions. To eliminate potential human bias, all rendered videos are generated using the fixed body shape. All segments are approximately 6 to 8 seconds long and should not be shorter than 5 seconds. They are complete phrases, beginning and ending at word boundaries, and do not end abruptly or on a “cliffhanger”. We recruit 40 participants from our lab [^1] https://april.zju.edu.cn/, all of whom have strong English proficiency and are ensured to be with no prior knowledge/awareness of this project before the user study. The majority are male (about 75%), with the remaining being female. Almost all participants are residents of mainland China, except for one international student from South Korea. Their ages range from 20 to 30 years old. Each participant receive a 9 GBP compensation, determined based on the average completion time (about 45 minutes) of our user study (including Sec. 4.2.3, Sec. 4.3.3, and Sec. 4.4.3) and the average wage level [71].
All tests are conducted using pairwise comparisons. Participants are shown two video clips synthesized by different models on each page. They are then required to label their preferred clip based on the instructions displayed below the videos. Besides, each participant is required to complete a substantial training page featuring a fixed set of videos before starting each test, to familiarize them with the task. Moreover, an attention check is randomly inserted into the experiment to ensure valid responses. Concretely, a text message reading: “Attention: Please select the right motion” appear both at the bottom of the video pair for the entire duration of the question and during the transition gap between the two clips. Any samples that fail the attention check are excluded from the final analysis. Another filtering is also employed, users are asked to provide a brief qualitative reason for their choice in a random page of each test. This written feedback helps us assess the reliability of the results.
After the above preliminaries, we follow the evaluation in GENEA [71] and propose three types of preference tests: one test assesses human-likeness, while the other two evaluate appropriateness in different aspects: speech-motion synchronization, face-body coordination. We report pairwise wining proportion in the human-likeness test and matched percentage in the appropriateness test. We adopt the pre-trained Combo on neutral data in these evaluations.
Figure 6: Five screenshots of the user interface from our studies. Please zoom in for more details.
Table III: Average scores of user study with 95% confidence intervals. In each test, GT serves as the topline, with higher scores indicating better performance across all metrics. We omit the %. Among all tests, TalkSHOW and EMAGE involve a single model, whereas our method involves both pretraining and finetuning versions.
Human-likeness Synchronized Coordinated Generalized-E Generalized-I
@@
[…]
Human-likeness. In the human-likeness test, participants are asked to evaluate “Please indicate which motion most closely resembles that of a real human”. For this test, we select 24 speech segments from the BEAT dataset’s test set of Speaker-2 (BEAT-S2 is chosen due to baseline models availability for this identity) to generate corresponding motions, resulting in 24 video clips for each method. These video clips are muted to eliminate any influence from the speech content. Four approaches yields 12 possible pairwise combinations for side-by-side comparisons. Each participant assess 24 video pairs, covering all 24 speech samples, with each of the 12 combinations appearing twice. The speech samples and the pairing of comparisons are randomized for each participant.
In this experiment, 40 participants pass the filtering checks, and the results are presented in Tab. III column 1. Our method outperforms the two recent baselines, which verifies that our method significantly surpasses the compared SOTA methods of human-likeness.
Synchronization. In the synchronization test, participants are asked to “Please indicate which motion best matches the speech, both in terms of rhythm and semantic”. We follow mismatching methodology as GENEA [71]. Specifically, we select 10 speech segments from BEAT-S2 to create 10 video clips per method. We then rearrange the speech-motion segments so that each motion is paired with an unsynchronized speech. This creates a pair of motion videos features the same speech audio, one matched and one mismatched. Notably, mismatched pairs may have misaligned durations. We address this using the method used in GENEA. The position of the mismatched video (left or right) is randomized. Each participant is asked to assess 40 pairs (10 speech segments $\times$ 4 approaches).
In this experiment, 37 subjects pass the filtering checks, and the results are shown in Tab. III column 2. Matched percentage identities how often subjects prefer matched over mismatched motion. Because the higher the motion-speech synchronization in the generated results, the more easily participants can detect mismatches. Consequently, our method is more synchronized for the specific speech signal compared to other recent methods.
Coordination. The coordination test is largely consistent with the synchronization test, with the only difference lying in the construction of the mismatched pairs. Participants are asked to ”Please indicate which motion has the better coordination between facial and body movements”. They are shown a pair of videos: one shows facial and body movements generated from the same speech segment, and the other are from different samples. The videos are muted in this test. In this experiment, 39 subjects pass the filtering checks and the results are shown in Tab. III column 3. The mismatched percentage shows how often subjects notice that the face and body motions come from different samples. The more harmonious the face and body motions are, the higher the rate at which mismatches are identified after being shuffled. The results verify that the body and face movements generated by our method are better aligned with each other due to the divide-and-unite design.
4.3 Verification on X-Adapter
In Sec. 4.3.1, we first analyze the effectiveness of the X-adapter in emotion transfer during training, and then further explore the two benefits it brings: multi-modal editing and unseen emotion editing during inference. These experiments are conducted on Scott data. Then, in Sec. 4.3.2, we analyze the effectiveness of the X-adapter in identity transfer and validate cross-identity transfer under various conditions. These experiments are conducted on Neutral data. We further conduct the user study to supplement more comprehensive subjective analysis in Sec. 4.3.3. Finally, we provide a detailed analysis of overall tuning efficiency in Sec. 4.3.4.